kasey0514:

If a 10-volt battery is placed across a 2-ohm resistance (R = 2Ω), a current of amperes will flow

Laylalyssa:

So basically, when solving voltage you use this formula V=IR in this V represents the voltage, where "I" represent the current and R represents the resistance. Then, the given values are enumerated below and these will help us in solving for the unknown. V=10 volts I=? R=2 ohms V=IR I=V/R I=10/2 I=5 amperes knowing this what do u think the answer is?